OpenStudy (anonymous):

Indicate the equation of the line meeting the given conditions. Please put the equation in standard form. Containing A(1, 3) and B(0, 2)

OpenStudy (anonymous):

Please explain how you did it too...

OpenStudy (kropot72):

Do you know the formula for finding the slope of the line?

OpenStudy (kropot72):

\[y-y _{1}=m(x-x _{1})\] where m is the slope. Can you put in the given values and solve to find m?

OpenStudy (kropot72):

(3 - 2) = m(1 - 0) m = 1 The standard form for the equation of a straight line is y = mx + b where b is the intercept on the y axis ( where x = 0). So in this case we know that b = 2. So our equation is y = x + 2

OpenStudy (kropot72):

@natnatwebb Do you understand how the equation was found?

OpenStudy (anonymous):

Not really, doesn't standard form have to be like y+x=whatever? Also, isn't the slope formula \[x_{2}-x _{1}/y_{2}-y _{1}\]

OpenStudy (anonymous):

you have it upside down \[m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

you are comparing the y value to the x value, like miles per gallon or price per ounce y's go up top, x's in the bottom

OpenStudy (anonymous):

so thats m, so you put in that number for m, put the y and x numbers in for x1 and y1, multiply the numbers, the just manipulate the equation till you have y and x on one side and a whole number on the other?

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

first we need the slope. we do it in our heads and then compute from \((0,2)\) to \((1,3)\) x increases by 1 unit (from 0 to 1) and so does y (from 2 to 3) so the slope is 1 now lets do it by computing \[m=\frac{3-2}{1-0}=\frac{1}{1}=1\]

OpenStudy (anonymous):

now that we have the slope, we need the \(y\) intercept, which is what you get when \(x=0\) from the point \((0,2)\) we know if \(x=0\) then \(y=2\) so the \(y\) intercept is 2

OpenStudy (anonymous):

we then get the equation right away \[y=1\times x+2\] but don't write \(1\times x\) because you teacher will think you are daft, so just write \[y=x+2\]

OpenStudy (anonymous):